OpenStudy (anonymous):

Given the following function, find: (a) vertex, (b) axis of symmetry, (c) intercepts, (d) domain, (e) range, (f) intervals where the function is increasing, (g) intervals where the function is decreasing, and (h) the graph of the function. Please show all of your work. f(x) = -x^2-6

OpenStudy (anonymous):

a) Vertex (0,-6) b)x=0 c)None. Two, but they are imaginary. d) Domain: \[(-\infty, \infty)\] e) Range: \[(-\infty, -6)\] f) Increasing interval \[(-\infty, 0)\] g) Decreasing interval \[(0, \infty)\] The graph is a parabola pointing down where the y coordinate is -6.

OpenStudy (anonymous):

what does this graph look like?!

OpenStudy (anonymous):



OpenStudy (anonymous):

-6, sorry

OpenStudy (anonymous):

great!~ Thank you so much!! :)